OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. x2 - 20x + 100 = 36 A. {-6, 6} B. {-5/3, 5/3} C. {6, 10} D. {4, 16} First factor the trinomial as a binomial square and then use the square root method to solve. x2 - 20x + 100 = 36 A. {-6, 6} B. {-5/3, 5/3} C. {6, 10} D. {4, 16} @Mathematics

hero (hero):

x^2-20x+64 = 0 x^2-16x-4x + 64 =0 x(x-16)-4(x-16)=0 (x-16)(x-4)=0 x =4,16